# WiLoR: End-to-end 3D Hand Localization and Reconstruction in-the-wild

### v1

@@ Abstract
Figure 1: We propose WiLoR, a full-stack in-the-Wild Localization and 3D hand Reconstruction method. WiLoR first localizes and defines the handedness of the detected hands which are then lifted to 3D using a transformer-based hand pose estimation module. To aid high-fidelity reconstructions and facilitate image-alignment, we employ a refinement module that extracts localized features to correct miss-aligned poses. WiLoR achieves state-of-the-art performance under different benchmark datasets while boosting the temporal coherence of image-based 3D hand pose estimation methods.
@@ 1 Introduction
Hand detection and reconstruction has been a long-studied problem due to its numerous applications, ranging from virtual reality [22] to sign language [2, 30] and human behaviour recognition [65]. Given the large variations in hand appearance and articulation [64] along with heavy occlusions and motion blur [24, 26] that are usually present in hand interactions, the task of hand pose estimation is a considerably challenging. Over the years, several methods have been proposed to tackle 3D hand pose estimation [49, 11, 42, 61]. However, despite producing credible results, these methods primarily focus on images containing a fixed number of hands and hence cannot generalize to in-the-wild images.
@@
In the closely related fields of 3D human body and face reconstruction, state-of-the-art methods [7, 17, 23, 21] employ bottom-up pipelines founded on top of high-performance detection models that initially localize human body and face within the image, enabling their generalization to in-the-wild images. Despite the numerous methods that have been proposed to solve the task of human body and face detection, there has been a notable lack in real-time hand detection methods. The importance of hand detectors is further emphasized considering that current 3D hand pose estimation frameworks operate on tight crops around the hand regions [10, 96], high fidelity detectors are essential for the generalisation of such methods in in-the-wild scenarios. Popular hand detection and localisation methods [92, 7], fail significantly to detect multiple hands and challenging poses, while more recent methods [52, 53, 39] albeit producing reasonable results can not operate in real-time. Motivated from the lack of accurate hand detection frameworks, we propose a robust single-state anchor-free detector that can operates in over than 100 frames-per-second (fps). As we experimentally showcase, robust detections can enforce more stable 4D reconstructions and overcome jittering artifacts which is currently one of the main limitations of 3D frame-based pose estimation methods.
@@
In contrast to the relatively unexplored hand detection and localization, 3D hand pose estimation has received significantly more attention. Initial 3D pose reconstruction methods have focused on traditional convolution-based backbones to process and extract image features [5, 36, 49, 11]. Following the success of transformers and their ability to consume large amounts of data [85, 14, 37, 70, 83], several methods have paved the way of utilising transformer architectures scaling up the 3D human body and hand recovery [41, 42, 61]. Recently, Pavlakos et al. [61] showcased the effectiveness of vision transformers (ViT) using a simple yet powerful framework trained on a large-scale dataset. The key to the success of this method lies in the scale of its architecture, which is composed of more than 0.5 billion parameters, enabling it to effectively consume large amount of data. However, as shown in the literature [55, 79, 40, 93], regressing the hand parameters from a single image results in bad alignment and incorrect poses. Currently, methods that aim to achieve better image alignment rely on a sub-optimal solutions, such as intermediate heatmap representations [34, 10, 96]. To tackle this, we propose a high-fidelity 3D pose estimation method that decomposes 3D hand reconstruction into two stages. In particular, the decoder first predicts a rough hand estimation that is used to extract multi-scale image-aligned features from our refinement module. By leveraging the rough hand estimation, we can extract meaningful spatial features that lead to better image alignment and state-of-the-art performance on FreiHand [100] and HO3D [24] benchmark datasets. Additionally, in contrast to vertex regression methods [36, 41, 42] that directly regress 3D vertices, our method predicts MANO parameters [71], ensuring both explainable and plausible hand poses.
@@
  • Based on the limitations of current hand detection benchmark datasets, we collect a large-scale dataset of in-the-wild images that contain multiple hands and introduce a challenging benchmark for hand detection. We make the dataset along with its 2D and 3D publicly available.
  • We propose a real-time hand detection method trained on the aforementioned large-scale dataset that outperforms previous hand detection methods by a large margin in both accuracy and efficiency.
  • We propose a transformer-based method that facilitates high fidelity 3D reconstructions that tackles the architectural limitations of previous method using an novel refinement module. The proposed method, apart from highly efficient, achieves state-of-the-art performance in Freihand and HO3D benchmark datasets.
@@ 2 Related Work
Hand Detection and Tracking. Object detection has been extensively studied in the literature achieving remarkable advancements [47, 44, 67] and setting the foundations for human body [80, 63, 19, 58, 17, 90] and face detection [12, 98] pipelines. In contrast, despite two decades of research efforts [69], hand detection has not yet achieved comparable breakthroughs. Initial approaches used controlled conditions and depth cameras [88, 97, 78, 87, 77] to detect and track human hands. Several efforts have been made to boost hand detection under different skin tones and backgrounds using multi-stage frameworks [48, 62], however, they fail to generalize in challenging environments. Following the success in object detection, several methods have adopted fully convolutional architectures for hand detection [29, 72, 13, 92, 73]. Simon et al. [75] introduced a multi-view bootstrapping procedure to annotate in-the-wild data and train a real-time convolutional detector network. Recently, Narasimhaswamy et al. [52] proposed an extension of MaskRCNN [28] network to detect in-the-wild hands and identifying their corresponding contact points [53] and body associations [54]. Nevertheless, despite the extensive efforts in the literature, most methods rely on slow backbones and struggle with challenging images. The primary issue is the lack of large-scale training data featuring multiple levels of occlusions and motion blur from in-the-wild scenes. To tackle such limitation, we propose lightweight hand detector that is 45 $\times$ faster compared to previous state-of-the-art detectors, trained on 2M in-the-wild images with diverse environments and occlusion.
@@
3D Hand Pose Estimation. Similar to hand detection, initial approaches for hand pose estimation relied on depth cameras [57, 81, 18] to reconstruct 3D hands. Boukhayma et al. [5] introduced the first fully learnable pipeline that directly regresses the parameters of the MANO hand model [71] from RGB images. In a similar manner, several follow-up works used heatmaps [94] and iterative refinement [1] to enforce 2D alignment. Kulon et al. [35, 36] introduced an alternative regression method that regresses 3D vertices instead of MANO pose parameters, which significantly outperformed previous methods. Various approaches have been proposed to improve task-specific challenges of 3D pose estimation, including robustness to occlusions [59] and motion blur [56] and reducing inference speed [10, 96]. Recently, Pavlakos et al. [61] highlighted the importance of scaling up both the training data and the capacity of the model. Specifically, building on the success of the Vision Transformer (ViT) backbones for body pose estimation [40, 20, 6], they demonstrated that using a simple yet effective large-scale transformer architecture can achieve state-of-the-art performance when trained on a diverse collection of datasets. However, directly regressing MANO parameters from the image in one go may introduce miss-alignments and incorrect poses. To tackle this, we propose a novel refinement layer that deforms hand pose using mesh-aligned multi-scale features.
@@ 3 WHIM Dataset
A vital cause behind the lack of high-fidelity hand detection systems lies in the limited amount of in-the-wild datasets with multiple hand annotations. To build a robust hand detection and reconstruction framework, we collected a large-scale dataset with millions of in-the-wild hands (WHIM) with diverse poses, illuminations, occlusions, and skin tones.
To collect the proposed dataset, we devised a pipeline to automatically annotate YouTube videos from diverse and challenging in-the-wild scenarios. In particular, we selected more than 1,500 YouTube videos containing hand activities including sign language, cooking, everyday activities, sports, and games with ego- and exo-centric viewpoints, motion blur, different hand scales, and interactions. To accurately detect and annotate the hands on each frame we used a combination of ensemble networks. Firstly, we used VitPose [90] and AlphaPose [17] to detect all humans in the frame and selected the bounding boxes with confidence bigger than 0.65. We then cropped the bounding boxes and fed them to an ensemble hand detection pipeline that consists of MediaPipe [92], OpenPose [7] and ContactHands [52] models. To localize the hand, we used a weighted average between the bounding box positions $\mathbf{b}_{i}$ of the three detectors $d_{i}$, scaled from their corresponding confidence $P(y_{i}|d_{i})$:
$\hat{y}=\frac{\sum_{i}P(\mathbf{b}_{i}|d_{i})y_{i}}{\sum_{i}P(\mathbf{b}_{i}|d_{i})}$ (1)
@@
where $\mathcal{L}_{BL}$ and $\mathcal{L}_{A}$ denote the bone length and the joint angle loss terms, respectively. For additional details of the bio-mechanical constraints, we refer the reader to [76].
Finally, given that the bio-mechanical prior acts mainly on the joint space, we followed [2] and trained a PCA model on ARCTIC dataset [16] acting as a 3D prior and to model the distribution of feasible hand poses. We formulated the prior loss as the reconstruction error of the 3D mesh $\mathbf{X}$ projected and reconstructed from the PCA space $\mathbf{U}$ as:
@@
where $\mathbf{U}\in\mathbb{R}^{N\cdot 3\times d}$ is the eigenvector basis of $d$ components and $\boldsymbol{\mu}$ is the mean mesh. In Fig. 2
@@ 4.1 Hand Detection and Localization
Over the past years, fully convolutional networks (FCNs) have shown remarkable efficiency in human detection [86] and object detection [68]. Building on their success, we employ an FCN architecture to achieve both accurate and real-time hand localization. Similar to object detection frameworks, given an image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ our goal is to detect the bounding boxes $\mathbf{B}=\{\mathbf{b}_{j}\in\mathbb{R}^{4}:0\leq j\leq n\}$ of the $n$ hands present in the image along with their hand side label $\mathbf{y}_{j}$. We follow the commonly used one-stage backbone-neck-head formulation, where we built upon the powerful and efficient DarkNet backbone [67]. We extract the last three feature maps $\{C3,C4,C5\}$ of the backbone to generate a multi-scale feature pyramid in the neck module. To enable our model to effectively capture multi-scale features using both top-down and bottom-up pathways across different feature maps, we utilized Path Aggregation Network (PANet) [45], an extension of Feature Pyramid Network [43] that facilitates fine-grained information flow using a bottom-up path augmentation. Finally, we use three detection heads to predict the bounding boxes $\mathbf{b}_{j}$ and hand side labels $\mathbf{y}_{j}$ at different anchor resolutions. Following [15], we adopt an anchor-free design to enhance the flexibility of our localization method and directly predict bounding box coordinates without relying on predefined anchor boxes. An overview of the proposed detection network is visualized in Fig. 3. Similar to [12], we observed that joint keypoint supervision significantly improved the performance and the robustness of the detector. The full training objective can be defined as:
@@
where $\mathcal{L}_{BCE}$ is the binary cross entropy loss between the predicted and the ground truth box labels, $\mathcal{L}_{DFL}$ denotes the distributional focal loss [38] which measures the difference between the predicted and the ground truth bounding box distributions, $\mathcal{L}_{CIoU}$ measures the discrepancy between the predicted and the ground truth bounding box [95] and $\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3}$ are weights that balance the losses. For additional details about the detection network training, we refer the reader to the supplementary material.
@@ 4.2 Hand Reconstruction
Given an image $\mathbf{I_{h}}\in\mathbb{R}^{H\times W\times 3}$ that contains a human hand, tightly cropped around the hand detectors bounding box, the proposed 3D hand reconstruction method estimates the corresponding hand pose $\theta\in\mathbb{R}^{45}$ and shape $\beta\in\mathbb{R}^{10}$ MANO [71] parameters along with the camera $\mathbf{K}_{cam}=\{\mathbf{t}_{cam},\mathbf{s}_{cam}\}$ to obtain a 3D hand.
To build a powerful 3D pose estimation network that can scale on large amounts of data, we follow [40, 6, 61], and built our backbone using a pre-trained ViT encoder [85, 90]. The image $\mathbf{I_{h}}$ is first split into $M$-size patches $\mathbf{P}\in\mathbb{R}^{\frac{HW}{M^{2}}\times(M^{2}\times 3)}$ and then embedded to high dimensional tokens $\mathbf{T}_{img}\in\mathbb{R}^{\frac{HW}{M^{2}}\times C}$. To uniquely encode their spatial location, positional embeddings $\mathbf{P}_{e}$ are added to the image tokens $\mathbf{T}_{img}$ [85]. In addition to the image tokens, we explicitly model hand pose, shape, and camera parameters with three distinct tokens $\mathbf{T}_{pose},\mathbf{T}_{shape},\mathbf{T}_{cam}$. We then feed the concatenated tokens to a ViT transformer encoder to obtain a set of updated feature tokens $\mathbf{T^{\prime}}_{img},\mathbf{T^{\prime}}_{pose},\mathbf{T^{\prime}}_{shape},\mathbf{T^{\prime}}_{cam}$. Using a set of MLP layers we regress a rough estimation of pose $\theta^{c}$ and shape $\beta^{c}$ parameters of the MANO model, which will serve as a prior for the refinement network. Similarly, we regress the camera translation and scale parameters $\mathbf{K}_{cam}=\{\mathbf{t}_{cam},\mathbf{s}_{cam}\}$ from the camera token features.
Multi-Scale Pose Refinement Module. In order to get better image alignment and more accurate hand pose, we introduce a fully differentiable refinement module that predicts pose and shape residuals of the rough hand estimation. To do so we leverage the image features extracted from the ViT backbone to serve as the 2D feature cues in our refinement module. In particular, we reshape image feature tokens $\mathbf{T^{\prime}}_{img}$ to form a low resolution feature map $\mathbf{F}_{0}\in\mathbb{R}^{\frac{H}{M}\times\frac{W}{M}\times C}$ and project the rough hand estimation $\mathcal{M}_{l}$ to feature map using the estimated $\mathbf{t}_{cam},\mathbf{s}_{cam}$ camera parameters. Then, using bilinear interpolation we sample from $\mathbf{F}_{0}$ a feature vector $\mathbf{f}^{\mathbf{v}}_{0}$ for each projected vertex $\mathbf{v}$:
@@
Given that the initial feature map is very low-dimensional, we use a set of deconvolutional layers to upsample $\mathbf{F}_{0}$ to multiple higher resolution feature maps ${\mathbf{F}_{0},\mathbf{F}_{1},...,\mathbf{F}_{n}}$ that will serve as multi-scale features for the proposed refinement module. Intuitively, low-dimensional feature maps will provide global and structural residuals of the hand shape while more high-resolution features could provide finer details of the hand pose.
Loss function. The method proposed in this paper is trained with supervision for 3D vertices $\hat{\mathbf{V}}_{3D}$, 2D joints $\hat{\mathbf{J}}_{2D}$ and MANO parameters $\hat{\theta},\hat{\beta}$, when available. Additionally, following [33, 61] we utilize a discriminator network $D$ to enforce plausible hand poses and shapes and penalize irregular articulations. The full loss function can be defined as:

### v2

@@ Abstract
Figure 1: We propose WiLoR, a full-stack in-the-Wild Localization and 3D hand Reconstruction method. WiLoR first localizes and defines the handedness of the detected hands which are then lifted to 3D using a transformer-based hand pose estimation module. To aid high-fidelity reconstructions and facilitate image-alignment, we introduce a refinement module that extracts localized features to correct misaligned poses. WiLoR achieves state-of-the-art performance under different benchmark datasets while boosting the temporal coherence of image-based 3D hand pose estimation methods.
@@ 1 Introduction
Hand detection and reconstruction has been a long-studied problem due to its numerous applications, ranging from virtual reality [23] to sign language [2, 108, 31] and human behaviour recognition [68]. Given the large variations in hand appearance and articulation [67] along with heavy occlusions and motion blur [25, 27] that are usually present in hand interactions, the task of hand pose estimation is considerably challenging. Over the years, several methods have been proposed to tackle 3D hand pose estimation [51, 12, 44, 63]. However, despite producing credible results, these methods primarily focus on images containing a fixed number of hands and hence cannot generalize to in-the-wild images.
@@
In the closely related fields of 3D human body and face reconstruction, state-of-the-art methods [8, 18, 24, 22] employ bottom-up pipelines founded on top of high-performance detection models that initially localize human body and face within the image, enabling their generalization to in-the-wild images. Despite the numerous methods that have been proposed to solve the task of human body and face detection, there has been a notable lack in real-time hand detection methods. The importance of hand detectors is further emphasized considering that current 3D hand pose estimation frameworks operate on tight crops around the hand regions [11, 103]. Consequently, hand detectors are essential for the generalisation of such methods in in-the-wild scenarios. Popular hand detection and localisation methods [97, 8] fail significantly to detect multiple hands and challenging poses, while more recent methods [54, 55, 41] albeit producing reasonable results struggle to operate in real-time. Motivated by the lack of accurate hand detection frameworks, we propose a robust single-state anchor-free detector that can operate in over than 100 frames-per-second (fps). As we experimentally show, robust detections can enforce more stable 4D reconstructions and overcome jittering artifacts which is currently one of the main limitations of 3D frame-based pose estimation methods.
@@
In contrast to the relatively unexplored hand detection and localization problem, 3D hand pose estimation has received significantly more attention. Initial 3D pose reconstruction methods have focused on traditional convolution-based backbones to process and extract image features [5, 38, 51, 12]. Following the success of transformers and their ability to consume large amounts of data [88, 15, 39, 73, 86, 99], several methods have paved the way of utilising transformer architectures scaling up the 3D human body and hand recovery [43, 44, 63]. Recently, Pavlakos et al. [63] showcased the effectiveness of vision transformers (ViT) using a simple yet powerful framework trained on a large-scale dataset. The key to the success of this method lies in the scale of its architecture, composed of more than 0.5 billion parameters, enabling effective consumption of large amount of data. However, as shown in the literature [57, 82, 42, 98], regressing the hand parameters from a single image results in poor alignment and incorrect poses. Currently, methods that aim to achieve better image alignment rely on sub-optimal solutions, such as intermediate heatmap representations [36, 11, 103]. To tackle this, we propose a high-fidelity 3D pose estimation method that decomposes 3D hand reconstruction into two stages. In particular, the decoder first predicts a rough hand estimation that is used to extract multi-scale image-aligned features from our refinement module. By leveraging the rough hand estimation, we can extract meaningful spatial features that lead to better image alignment and state-of-the-art performance on FreiHand [107] and HO3D [25] benchmark datasets. Additionally, in contrast to vertex regression methods [38, 43, 44] that directly regress 3D vertices, our method predicts MANO parameters [74], ensuring both explainable and plausible hand poses.
@@
  • Based on the limitations of current hand detection benchmark datasets, we collect a large-scale dataset of in-the-wild images that contain multiple hands and introduce a challenging benchmark for hand detection. We make the dataset, along with the corresponding 2D and 3D annotations, publicly available.
  • We propose a real-time hand detection method trained on the collected large-scale dataset that outperforms previous hand detection methods by a large margin in both accuracy and efficiency.
  • We propose a transformer-based method that facilitates high fidelity 3D reconstructions and tackles the architectural limitations of previous methods using a novel refinement module. The proposed method, apart from highly efficient, achieves state-of-the-art performance in both Freihand and HO3D benchmark datasets.
@@ 2 Related Work
Hand Detection and Tracking. Object detection has been extensively studied in the literature achieving remarkable advancements [49, 46, 70] and setting the foundations for human body [83, 65, 20, 60, 18, 94] and face detection [13, 105] pipelines. In contrast, despite two decades of research efforts [72], hand detection has not yet achieved comparable breakthroughs. Initial approaches used controlled conditions and depth cameras [91, 104, 81, 90, 80] to detect and track human hands. Several efforts have been made to boost hand detection under different skin tones and backgrounds using multi-stage frameworks [50, 64], however, they fail to generalize in challenging environments. Following the success in object detection, several methods have adopted fully convolutional architectures for hand detection [30, 75, 14, 97, 76]. Simon et al. [78] introduced a multi-view bootstrapping procedure to annotate in-the-wild data and train a real-time convolutional detector network. Recently, Narasimhaswamy et al. [54] proposed an extension of MaskRCNN [29] network to detect in-the-wild hands and identifying their corresponding contact points [55] and body associations [56]. Nevertheless, despite the extensive efforts in the literature, most methods rely on slow backbones and struggle with challenging images. The primary issue is the lack of large-scale training data featuring multiple levels of occlusions and motion blur from in-the-wild scenes. To tackle such limitation, we propose a lightweight hand detector that is 45 $\times$ faster compared to previous state-of-the-art detectors, trained on 2M in-the-wild images with diverse environments and occlusion.
@@
3D Hand Pose Estimation. Similar to hand detection, initial approaches for hand pose estimation relied on depth cameras [59, 84, 19] to reconstruct 3D hands. Boukhayma et al. [5] introduced the first fully learnable pipeline that directly regresses the parameters of the MANO hand model [74] from RGB images. In a similar manner, several follow-up works used heatmaps [100] and iterative refinement [1] to enforce 2D alignment. Kulon et al. [37, 38] introduced an alternative regression method that directly regresses 3D vertices using spiral graph neural networks [6, 66], which significantly outperformed previous methods. Various approaches have been proposed to improve task-specific challenges of 3D pose estimation, including robustness to occlusions [61] and motion blur [58] and reducing inference speed [11, 103]. Recently, Pavlakos et al. [63] highlighted the importance of scaling up both the training data and the capacity of the model. Specifically, building on the success of the Vision Transformer (ViT) backbones for body pose estimation [42, 21, 7], they demonstrated that using a simple yet effective large-scale transformer architecture can achieve state-of-the-art performance when trained on a diverse collection of datasets. However, directly regressing MANO parameters from the image in one go may introduce misalignments and incorrect poses. To tackle this, we propose a novel refinement layer that deforms hand pose using mesh-aligned multi-scale features.
@@ 3 WHIM Dataset
A vital cause behind the lack of high-fidelity hand detection systems lies in the limited amount of in-the-wild datasets with multiple hand annotations. To build a robust hand detection framework, we collected a large-scale dataset with millions of in-the-wild hands (WHIM) with diverse poses, illuminations, occlusions, and skin tones.
To collect the proposed dataset, we devised a pipeline to automatically annotate YouTube videos from diverse and challenging in-the-wild scenarios. In particular, we selected more than 1,400 YouTube videos containing hand activities including sign language, cooking, everyday activities, sports, and games with ego- and exo-centric viewpoints, motion blur, different hand scales, and interactions. To accurately detect and annotate the hands on each frame we used a combination of ensemble networks. First, we used VitPose [94] and AlphaPose [18] to detect all humans in the frame and selected the bounding boxes with confidence bigger than 0.65. We then cropped the bounding boxes and fed them to an ensemble hand detection pipeline that consists of MediaPipe [97], OpenPose [8] and ContactHands [54] models. To localize the hand, we used a weighted average between the bounding box positions $\mathbf{b}_{i}$ of the three detectors $d_{i}$, scaled from their corresponding confidence $P(\mathbf{b}_{i}|d_{i})$:
$\hat{y}=\frac{\sum_{i}P(\mathbf{b}_{i}|d_{i})\mathbf{b}_{i}}{\sum_{i}P(\mathbf{b}_{i}|d_{i})}$ (1)
@@
where $\mathcal{L}_{BL}$ and $\mathcal{L}_{A}$ denote bone length and joint angle loss terms, respectively. For additional details of the bio-mechanical constraints we refer the reader to [79].
Finally, given that the bio-mechanical prior acts mainly on the joint space, we followed [2] and trained a PCA model on ARCTIC dataset [17] acting as a 3D prior, modeling the distribution of feasible hand poses. We formulated the prior loss as the reconstruction error of the 3D mesh $\mathbf{X}$ projected and reconstructed from the PCA space $\mathbf{U}$, as:
@@
where $\mathbf{U}\in\mathbb{R}^{d\times N\cdot 3}$ denotes the eigenvector basis of $d$ components and $\boldsymbol{\mu}$ the mean mesh. Fig. 2 includes several examples of the proposed WHIM dataset.
@@ 4.1 Hand Detection and Localization
Over the past years, fully convolutional networks (FCNs) have shown remarkable efficiency in human detection [89] and object detection [71]. Building on their success, we employ an FCN architecture to achieve both accurate and real-time hand localization. Similar to object detection frameworks, given an image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ our goal is to detect the bounding boxes $\mathbf{B}=\{\mathbf{b}_{j}\in\mathbb{R}^{4}:0\leq j\leq n\}$ of the $n$ hands present in the image along with their hand side label $\mathbf{y}_{j}$. We follow the commonly used one-stage backbone-neck-head formulation and we built upon the powerful and efficient DarkNet backbone [70]. We extract the last three feature maps $\{C3,C4,C5\}$ of the backbone to generate a multi-scale feature pyramid in the neck module. To enable our model to effectively capture multi-scale features using both top-down and bottom-up pathways across different feature maps, we utilized Path Aggregation Network (PANet) [47], an extension of Feature Pyramid Network [45] that facilitates fine-grained information flow using a bottom-up path augmentation. Finally, we use three detection heads to predict the bounding boxes $\mathbf{b}_{j}$ and hand side labels $\mathbf{y}_{j}$ at different anchor resolutions. Following [16], we adopt an anchor-free design to enhance the flexibility of our localization method and directly predict bounding box coordinates without relying on predefined anchor boxes. An overview of the proposed detection network is visualized in Fig. 3. Similar to [13], we observed that joint keypoint supervision significantly improved the performance and the robustness of the detector. The full training objective can be defined as:
@@
where $\mathcal{L}_{BCE}$ is the binary cross entropy loss between the predicted and the ground truth box labels, $\mathcal{L}_{DFL}$ denotes the distributional focal loss [40] which measures the difference between the predicted and the ground truth bounding box distributions, $\mathcal{L}_{CIoU}$ measures the discrepancy between the predicted and the ground truth bounding box [101], $\mathcal{L}_{kpts}$ denotes an $L2$ loss on the and $\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3}$ are weights that balance the losses.
@@ 4.2 Hand Reconstruction
Given an image $\mathbf{I_{h}}\in\mathbb{R}^{H\times W\times 3}$ that contains a human hand, tightly cropped around the hand detectors bounding box, the proposed 3D hand reconstruction method estimates the corresponding hand pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ MANO [74] parameters along with the camera parameters $\mathbf{K}_{cam}=\{\mathbf{t}_{cam},\mathbf{s}_{cam}\}$ to obtain a 3D hand.
To build a powerful 3D pose estimation network that can scale on large amounts of data, we follow [42, 7, 63], and built our backbone using a pre-trained ViT encoder [88, 94]. The image $\mathbf{I_{h}}$ is first split into $M$-size patches $\mathbf{P}\in\mathbb{R}^{\frac{HW}{M^{2}}\times(M^{2}\times 3)}$ and then embedded to high dimensional tokens $\mathbf{T}_{img}\in\mathbb{R}^{\frac{HW}{M^{2}}\times C}$. To uniquely encode their spatial location, positional embeddings $\mathbf{P}_{e}$ are added to the image tokens $\mathbf{T}_{img}$ [88]. In addition to the image tokens, we explicitly model hand pose, shape, and camera parameters with three distinct tokens $\mathbf{T}_{pose},\mathbf{T}_{shape},\mathbf{T}_{cam}$. We then feed the concatenated tokens to the ViT transformer encoder to obtain a set of updated feature tokens $\mathbf{T^{\prime}}_{img},\mathbf{T^{\prime}}_{pose},\mathbf{T^{\prime}}_{shape},\mathbf{T^{\prime}}_{cam}$. Using a set of MLP layers we regress a rough estimation of pose $\theta^{c}$ and shape $\beta^{c}$ parameters of the MANO model, which will serve as a prior for the refinement network. Similarly, we regress the camera translation and scale parameters $\mathbf{K}_{cam}=\{\mathbf{t}_{cam},\mathbf{s}_{cam}\}$ from the camera token features.
Multi-Scale Pose Refinement Module. In order to get better image alignment and more accurate hand pose, we introduce a fully differentiable refinement module that predicts pose and shape residuals of the rough hand estimation. To achieve this, we utilize image features extracted from the ViT backbone as 2D feature cues within our refinement module. In particular, we reshape image feature tokens $\mathbf{T^{\prime}}_{img}$ to form a low resolution feature map $\mathbf{F}_{0}\in\mathbb{R}^{\frac{H}{M}\times\frac{W}{M}\times C}$ and project the rough hand estimation $\mathcal{M}_{l}$ to feature map using the estimated $\mathbf{t}_{cam},\mathbf{s}_{cam}$ camera parameters. Then, using bilinear interpolation we sample from $\mathbf{F}_{0}$ a feature vector $\mathbf{f}^{\mathbf{v}}_{0}$ for each projected vertex $\mathbf{v}$:
@@
Given that the initial feature map is very low-dimensional, we use a set of deconvolutional layers to upsample $\mathbf{F}_{0}$ to multiple higher resolution feature maps ${\mathbf{F}_{0},\mathbf{F}_{1},...,\mathbf{F}_{n}}$ that will serve as multi-scale features for the proposed refinement module. Intuitively, low-dimensional feature maps will provide global and structural residuals of the hand shape while more high-resolution features provide finer details of the hand pose.
Loss function. The proposed model is trained with supervision for 3D vertices $\hat{\mathbf{V}}_{3D}$, 2D joints $\hat{\mathbf{J}}_{2D}$ as well as MANO parameters $\hat{\theta},\hat{\beta}$, when available. Additionally, following [35, 63], we utilize a discriminator network $D$ to enforce plausible hand poses and shapes and penalize irregular articulations. The full loss function can be defined as: